OpenStudy (anonymous):

Technetium-99m, has a half life of about 6 hours find the decay constant. Use the decay function to determine the amount of a 250mg dose that remains after 24 hours.

OpenStudy (anonymous):

I found the decay constant which is .1155 I don't know how to do the second part.

OpenStudy (wolf1728):

I don't know anything about the decay constant, but I know a LOT about half-life calculations. (He said modestly). LOL You can find a half life calculator here: http://www.1728.org/halflife.htm Technetium 99 half life about 6 hours you have 250 mg. How much after 24 hours? Here's the formula: Ending Amount = Beginning Amount / 2^n where n is the number of half lives the half life is about 6 hours so that means after 24 hours, four half lives have transpired. Ending Amount = 250 / 2^4 Ending Amount = 250 / 16 Ending Amount = 15.625 milligrams If you think that is incorrect, go the the calculator page here: http://www.1728.org/halflife.htm and check the calculations. You will get an answer of 15.625